# GC-MVSNet: Multi-View, Multi-Scale, Geometrically-Consistent Multi-View Stereo

### v3

@@ 3 Methodology
Our goal is to take $N$ views as input, including a reference image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and its paired $N$-1 source view images {$I_{i}$}${}^{N-1}_{i=1}$, along with the corresponding camera parameters $c_{0},...,c_{N}$, and then to estimate the reference view depth map ($D_{0}$) as the output.
@@ 3.2 Multi-View Geometric Consistency Module
GC-MVSNet estimates reference depth maps at three stages with different resolutions. At each stage, the GC module takes $D_{0}$, $M$ source view ground truths $D^{gt}_{1},...D^{gt}_{M}$, and their camera parameters $c_{0},...c_{M}$ as input (see Alg. 1). The GC module is then initialized with a geometric inconsistency mask sum (or $mask\_sum$) of zero at each stage. This mask sum accumulates the inconsistency of each pixel across the $M$ source views. For each source view, the GC module performs forward-backward reprojection of $D_{0}$ to generate the penalty and then adds it to the mask sum.
@@
Output: $per\_pixel\_penalty$
@@
    $D^{\prime\prime}_{P^{\prime\prime}_{0}},P^{\prime\prime}_{0}\leftarrow FBR(D_{0},c_{0},D_{i}^{gt},c_{i}^{gt})$triangleright Alg. 2
    $PDE\leftarrow{||P_{0}-P^{\prime\prime}_{0}||_{2}}$
    $RDD\leftarrow{\nicefrac{{1}}{{D_{0}}}||D^{\prime\prime}_{P^{\prime\prime}_{0}}-D_{0}||_{1}}$
@@
For each stage, we generate two binary masks of inconsistent pixels, $PDE_{mask}$ and $RDD_{mask}$, by applying thresholds $D_{pixel}$ and $D_{depth}$, and then take a logical-OR of the two to preduce a single mask of inconsistent pixels. These inconsistent pixels are assigned a value $1$ and all other pixels, including the consistent and the out-of-scope pixels, are assigned 00 to form a penalty mask. This penalty mask is then added to the mask sum (Alg. 1), which accumulates the penalty mask for each of the $M$ source views to generate a final mask sum with values $\in[0,M]$. Each pixel value indicates the number of inconsistencies of the pixel across the $M$ source views.
@@
$D_{(R\rightarrow S)}\leftarrow K_{S}\cdot E_{S}\cdot E_{R}^{-1}\cdot K_{R}^{-1}\cdot D_{0}$triangleright Project
@@
$D_{S_{remap}}\leftarrow REMAP(D_{i}^{gt},X_{D_{(R\rightarrow S)}},Y_{D_{(R\rightarrow S)}})$triangleright Remap
$D^{\prime\prime}_{P^{\prime\prime}_{0}}\leftarrow K_{R}\cdot E_{R}\cdot E_{S}^{-1}\cdot K_{S}^{-1}\cdot D_{S_{remap}}$triangleright Back project
@@ 3.3 Cost Function
$\begin{split}\small{L_{i}}_{(stage)}&=mean(\xi_{p}\odot\xi_{d})\\ \mathcal{L}_{total}&=\alpha.L_{1}+\beta.L_{2}+\gamma.L_{3}\end{split}$ (2)
@@ 4 Experiments
Learning-based
SurfaceNet [16] 0.450 1.040 0.745
@@
Table 1: Quantitative results on DTU evaluation set at 864 $\times$ 1152 resolution. Accuracy (Acc), completeness (comp) and overall are in $mm$. *** means that GBiNet is re-tested with the same post-processing threshold to all scans for fair comparison with other methods. Gray font shows the methods that use scan-specific thresholds for evaluation. Bold and underline represents first and second place, respectively.
@@ 4.1 Implementation Details
Following the general practice [6], we first train and evaluate our model on DTU. Then, we finetune on BlendedMVS to evaluate on Tanks and Temples. For training on DTU, we set the number of input images $N=5$ and image resolution as $512\times 640$. The depth hypotheses are sampled from $425mm$ to $935mm$ for coarse-to-fine regularization with the number of plane sweeping depth hypotheses for the three stages set to 48, 32, and 8. The corresponding depth interval ratio (DIR) is set as 2.0, 0.8, and 0.4. The model is trained with Adam [19] for 9 epochs with an initial learning rate ($LR_{DTU}$) of 0.001, which decays by a factor of 0.5 once after $8^{th}$ epoch. For the Geometric Consistency (GC) module, we use $M$=8 and set the stage-wise thresholds $D_{pixel}$ as 1, 0.5, 0.25 and ${D_{depth}}$ as 0.01, 0.005, 0.0025. We use $\alpha$=$\beta$=1 and $\gamma=2$ for all experiments. We train our model with a batch size of 3 on 8 NVIDIA RTX A6000 GPUs for about 9 hours.
@@
Method Mean normal-$\uparrow$ Fam. Fra. Hor. Lig. M60 Pan. Pla. Tra. Mean normal-$\uparrow$ Aud. Bal. Cour. Mus. Pal. Tem.
@@ 4.2 Experimental Performance
On DTU, we generate depth maps with $N$=5 at an input resolution of $864\times 1152$. We slightly adjust the depth interval ratio (DIR) to $1.6,0.7,0.3$ to accommodate the resolution change (more on DIR in Appendix C in Supplemental Material) and use the Fusibile algorithm [10] for depth fusion. Table 1 shows quantitative evaluations, where accuracy is the mean absolute distance in $mm$ from the reconstructed point cloud to the ground truth point cloud, completeness measures the opposite (see Appendix F in Supplemental Material), and overall is the average of these metrics, indicating the overall performance of the models. We find that GC-MVSNet achieves the best overall score as well as the best completeness score, when compared to nearly two dozen previous and state-of-the-art techniques. A qualitative evaluation is presented in Fig. 3 on a few sample MVS problems. We find that our model generates denser and more complete point clouds.
@@
Unlike DTU and Tanks and Temples, evaluation on Blended MVS is usually measured as the quality of depth maps, not the quality of point clouds. We set $N$=5, $M$=8, image resolution as $576\times 768$, and number of depth planes $D$=128, and finetune for 10 epochs with one-tenth the learning rate we used for DTU ($\frac{1}{10}LR_{DTU}$). We follow [5] for evaluation process.
@@
We also test the performance of our model on an outdoor dataset with the Tanks and Temples benchmark. To adapt to this change, we first finetune our model on BlendedMVS and then evaluate on the intermediate and advanced test sets of Tanks and Temples. We use an image resolution of $576\times 768$, $N$=7, $M$=10, one-tenth the learning rate of DTU ($\frac{1}{10}LR_{DTU}$), and $D$=192 for finetuning. We finetune the model for 12 epochs. The camera parameters and neighboring view selection are used as in R-MVSNet [49] and follow evaluation steps described in CDS-MVSNet [11].
@@ 4.3 Ablation Study
$\xi_{p}$ is generated using the mask sum ($mask\_sum$ in Alg. 1). It is the sum of penalties accumulated across the $M$ source views during multi-view geometric consistency check. At this stage, its elements take a discrete value between 00 and $M$. Using mask sum as it is leads to very high penalty per-pixel and consequently, very high loss value. Such a high loss value destabilizes the learning process. We control the magnitude of penalty by controlling the range of per-pixel penalty.
We explore two different ranges to control the magnitude of $\xi_{p}$, $[1,2]$ and $[1,3]$. To generate $\xi_{p}\in[1,2]$, we divide the mask sum by $M$ and then add $1$. To generate $\xi_{p}\in[1,3]$, we divide the mask by $\nicefrac{{M}}{{2}}$ and then add $1$. Table 5 shows the impact of these two ranges of penalty for $M$=8. Since $\xi_{p}\in[1,2]$ produces best results, we use it for all other experiments.
@@
The global hyperparameter $M$ is the number of source views across which the geometric consistency is checked, and is the same for all three stages (coarse, intermediate, and refinement stages). For training on DTU, we vary the value of $M$ while keeping $N$=5, i.e. while the MVS method uses only $4$ source views to estimate $D_{0}$, the GC module checks the geometrical consistency of $D_{0}$ across $M$ source views. It is important to note that the first $N$-1 out of $M$ source views are exactly the same used by GC-MVSNet to estimate $D_{0}$. We always keep $M\geq N-1$.
Table 5 presents a quantitative comparison for different values of $M$ and the amount of training iterations required for optimal performance of the model. At $M$=$N$-1= 4, i.e. checking geometric consistency across the same number of source views as used by GC-MVSNet to estimate $D_{0}$, the model performance significantly improves with sharp decrease in training iteration requirements, as compared to our baseline TransMVSNet-B (Table 7). As we increase the value of $M$ from $4$ to $10$, the training iteration required by our model further decreases. We find that at $M=8$, which is twice the number of source views used by GC-MVSNet, it achieves its best performance.
@@
Table 4: Impact of range of $\xi_{p}$ during training on DTU with $M$=8, $N$=5. Numbers are generated on DTU evaluation set.
@@
Table 4: Impact of range of $\xi_{p}$ during training on DTU with $M$=8, $N$=5. Numbers are generated on DTU evaluation set.
@@
Table 6: Overall score on the evaluation set of DTU [15] for different values of $D_{depth}$ and $D_{pixel}$. $M$ is fixed at $8$. C, I, and R means Coarse, Intermediate and Refine stages.
@@ C I R C I R
GC as a plug-in
CasMVSNet $L_{1}$$\times$ $\times$ 0.355 16
$L_{1}$checkmark $\times$ 0.357 16
@@
Stages
TransMVSNet-B CE $\times$ $\times$ 0.332 16
@@
Table 6: Overall score on the evaluation set of DTU [15] for different values of $D_{depth}$ and $D_{pixel}$. $M$ is fixed at $8$. C, I, and R means Coarse, Intermediate and Refine stages.

### v4

@@ 3 Methodology
Our goal is to take $N$ views as input, including a reference image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and its paired NN-1 source view images $I_{i}$$^{N-1}_{i=1}$, along with the corresponding camera parameters $c_{0},...,c_{N}$, and then to estimate the reference view depth map ($D_{0}$) as the output.
@@ 3.2 Multi-View Geometric Consistency Module
GC-MVSNet estimates reference depth maps at three stages with different resolutions. At each stage, the GC module takes $D_{0}$, $M$ source view ground truths $D^{gt}_{1},...D^{gt}_{M}$, and their camera parameters $c_{0},...c_{M}$ as input (see Alg. 1). The GC module is then initialized with a geometric inconsistency mask sum (or mask_summask\_sum) of zero at each stage. This mask sum accumulates the inconsistency of each pixel across the $M$ source views. For each source view, the GC module performs forward-backward reprojection of $D_{0}$ to generate the penalty and then adds it to the mask sum.
@@
Output: per_pixel_penaltyper\_pixel\_penalty
@@
  $D^{\prime\prime}_{P^{\prime\prime}_{0}},P^{\prime\prime}_{0}\leftarrow FBR(D_{0},c_{0},D_{i}^{gt},c_{i}^{gt})$ $\triangleright$ Alg. 2
[…]
@@
For each stage, we generate two binary masks of inconsistent pixels, $PDE_{mask}$ and $RDD_{mask}$, by applying thresholds $D_{pixel}$ and $D_{depth}$, and then take a logical-OR of the two to preduce a single mask of inconsistent pixels. These inconsistent pixels are assigned a value $1$ and all other pixels, including the consistent and the out-of-scope pixels, are assigned 0 to form a penalty mask. This penalty mask is then added to the mask sum (Alg. 1), which accumulates the penalty mask for each of the $M$ source views to generate a final mask sum with values $\in[0,M]$. Each pixel value indicates the number of inconsistencies of the pixel across the $M$ source views.
@@
$D_{(R\rightarrow S)}\leftarrow K_{S}\cdot E_{S}\cdot E_{R}^{-1}\cdot K_{R}^{-1}\cdot D_{0}$ $\triangleright$ Project
@@
$D_{S_{remap}}\leftarrow REMAP(D_{i}^{gt},X_{D_{(R\rightarrow S)}},Y_{D_{(R\rightarrow S)}})$ $\triangleright$ Remap
$D^{\prime\prime}_{P^{\prime\prime}_{0}}\leftarrow K_{R}\cdot E_{R}\cdot E_{S}^{-1}\cdot K_{S}^{-1}\cdot D_{S_{remap}}$ $\triangleright$ Back project
@@ 3.3 Cost Function
$\begin{split}{L_{i}}_{(stage)}&=mean(\xi_{p}\odot\xi_{d})\\ \mathcal{L}_{total}&=\alpha.L_{1}+\beta.L_{2}+\gamma.L_{3}\end{split}$ (2)
@@ 4 Experiments
Learning-based SurfaceNet [16] 0.450 1.040 0.745
@@
Table 1: Quantitative results on DTU evaluation set at 864 $\times$ 1152 resolution. Accuracy (Acc), completeness (comp) and overall are in $mm$. * means that GBiNet is re-tested with the same post-processing threshold to all scans for fair comparison with other methods. Gray font shows the methods that use scan-specific thresholds for evaluation. Bold and underline represents first and second place, respectively.
@@ 4.1 Implementation Details
Following the general practice [6], we first train and evaluate our model on DTU. Then, we finetune on BlendedMVS to evaluate on Tanks and Temples. For training on DTU, we set the number of input images $N=5$ and image resolution as $512\times 640$. The depth hypotheses are sampled from $425mm$ to $935mm$ for coarse-to-fine regularization with the number of plane sweeping depth hypotheses for the three stages set to 48, 32, and 8. The corresponding depth interval ratio (DIR) is set as 2.0, 0.8, and 0.4. The model is trained with Adam [19] for 9 epochs with an initial learning rate ($LR_{DTU}$) of 0.001, which decays by a factor of 0.5 once after $8^{th}$ epoch. For the Geometric Consistency (GC) module, we use MM=8 and set the stage-wise thresholds $D_{pixel}$ as 1, 0.5, 0.25 and $D_{depth}$} as 0.01, 0.005, 0.0025. We use $\alpha$$\beta$=1 and $\gamma=2$ for all experiments. We train our model with a batch size of 3 on 8 NVIDIA RTX A6000 GPUs for about 9 hours.
@@
Method Mean $\uparrow$ Fam. Fra. Hor. Lig. M60 Pan. Pla. Tra. Mean $\uparrow$ Aud. Bal. Cour. Mus. Pal. Tem.
@@ 4.2 Experimental Performance
On DTU, we generate depth maps with NN=5 at an input resolution of $864\times 1152$. We slightly adjust the depth interval ratio (DIR) to $1.6,0.7,0.3$ to accommodate the resolution change (more on DIR in Appendix C in Supplemental Material) and use the Fusibile algorithm [10] for depth fusion. Table 1 shows quantitative evaluations, where accuracy is the mean absolute distance in $mm$ from the reconstructed point cloud to the ground truth point cloud, completeness measures the opposite (see Appendix F in Supplemental Material), and overall is the average of these metrics, indicating the overall performance of the models. We find that GC-MVSNet achieves the best overall score as well as the best completeness score, when compared to nearly two dozen previous and state-of-the-art techniques. A qualitative evaluation is presented in Fig. 3 on a few sample MVS problems. We find that our model generates denser and more complete point clouds.
@@
Unlike DTU and Tanks and Temples, evaluation on Blended MVS is usually measured as the quality of depth maps, not the quality of point clouds. We set NN=5, MM=8, image resolution as $576\times 768$, and number of depth planes DD=128, and finetune for 10 epochs with one-tenth the learning rate we used for DTU ($\frac{1}{10}LR_{DTU}$). We follow [5] for evaluation process.
@@
We also test the performance of our model on an outdoor dataset with the Tanks and Temples benchmark. To adapt to this change, we first finetune our model on BlendedMVS and then evaluate on the intermediate and advanced test sets of Tanks and Temples. We use an image resolution of $576\times 768$, NN=7, MM=10, one-tenth the learning rate of DTU ($\frac{1}{10}LR_{DTU}$), and DD=192 for finetuning. We finetune the model for 12 epochs. The camera parameters and neighboring view selection are used as in R-MVSNet [49] and follow evaluation steps described in CDS-MVSNet [11].
@@ 4.3 Ablation Study
$\xi_{p}$ is generated using the mask sum (mask_summask\_sum in Alg. 1). It is the sum of penalties accumulated across the $M$ source views during multi-view geometric consistency check. At this stage, its elements take a discrete value between 0 and $M$. Using mask sum as it is leads to very high penalty per-pixel and consequently, very high loss value. Such a high loss value destabilizes the learning process. We control the magnitude of penalty by controlling the range of per-pixel penalty.
We explore two different ranges to control the magnitude of $\xi_{p}$, $[1,2]$ and $[1,3]$. To generate $\xi_{p}\in[1,2]$, we divide the mask sum by $M$ and then add $1$. To generate $\xi_{p}\in[1,3]$, we divide the mask by $\nicefrac{{M}}{{2}}$ and then add $1$. Table 5 shows the impact of these two ranges of penalty for MM=8. Since $\xi_{p}\in[1,2]$ produces best results, we use it for all other experiments.
@@
The global hyperparameter $M$ is the number of source views across which the geometric consistency is checked, and is the same for all three stages (coarse, intermediate, and refinement stages). For training on DTU, we vary the value of $M$ while keeping NN=5, i.e. while the MVS method uses only $4$ source views to estimate $D_{0}$, the GC module checks the geometrical consistency of $D_{0}$ across $M$ source views. It is important to note that the first NN-1 out of $M$ source views are exactly the same used by GC-MVSNet to estimate $D_{0}$. We always keep $M\geq N-1$.
Table 5 presents a quantitative comparison for different values of $M$ and the amount of training iterations required for optimal performance of the model. At MM=NN-1= 4, i.e. checking geometric consistency across the same number of source views as used by GC-MVSNet to estimate $D_{0}$, the model performance significantly improves with sharp decrease in training iteration requirements, as compared to our baseline TransMVSNet-B (Table 7). As we increase the value of $M$ from $4$ to $10$, the training iteration required by our model further decreases. We find that at $M=8$, which is twice the number of source views used by GC-MVSNet, it achieves its best performance.
@@
Table 4: Impact of range of $\xi_{p}$ during training on DTU with MM=8, NN=5. Numbers are generated on DTU evaluation set.
@@ C I R C I R
Table 6: Overall score on the evaluation set of DTU [15] for different values of $D_{depth}$ and $D_{pixel}$. $M$ is fixed at 88. C, I, and R means Coarse, Intermediate and Refine stages.
@@
GC as a plug-in CasMVSNet $L_{1}$ $\times$ $\times$ 0.355 16
$L_{1}$ $\checkmark$ $\times$ 0.357 16
@@
Stages TransMVSNet-B CE $\times$ $\times$ 0.332 16